OpenStudy (anonymous):

please help? Use the product-to-sum formula to write the given product as a sum or difference 12sin(pi/6)sin(pi/6)

OpenStudy (anonymous):

Do you have a specific product-to-sum formula? Or is it the one I'm thinking of? \[\cos(x\pm y)=\cos x\cos y\mp\sin x\sin y\\ \sin(x\pm y)=\sin x\cos y\pm \sin y\cos x\]

OpenStudy (anonymous):

it doesnt day a specific one

OpenStudy (jdoe0001):

I gather your case there uses the last indentity listed in there

OpenStudy (anonymous):

ANSWER CHOICES: 6sin(pi/12) 6-6cos(pi/3) 6+6cos(pi/12) -6sin(pi/12) 6sin(pi/6)+6cos(pi/6) if that helps

OpenStudy (jdoe0001):

so... what do you get using the trig identity there?

OpenStudy (anonymous):

I think its 6+6cos(pi/12)?

OpenStudy (jdoe0001):

I got something else :/

OpenStudy (anonymous):

you walk me through the steps? i did it using that and got a few different answers

OpenStudy (jdoe0001):

\(\bf 12\left[sin \left(\cfrac{\pi}{6}\right)sin \left(\cfrac{\pi}{6}\right)\right]\\ \left[sin \left(\cfrac{\pi}{6}\right)sin \left(\cfrac{\pi}{6}\right)\right] = \frac{1}{2}\left[ cos\left(\cfrac{\pi}{6}-\cfrac{\pi}{6}\right)-cos\left(\cfrac{\pi}{6}+\cfrac{\pi}{6}\right)\right]\)

OpenStudy (jdoe0001):

so notice the 1st pair subtracting from each other, what would \(\bf cos\left(\cfrac{\pi}{6}-\cfrac{\pi}{6}\right)\) be equal to?

OpenStudy (anonymous):

cos0?

OpenStudy (jdoe0001):

yes what about the 2nd pair, what would \(\bf cos\left(\cfrac{\pi}{6}+\cfrac{\pi}{6}\right)\) be equals to?

OpenStudy (anonymous):

sorry im confused.. im going guess 12(pi/6)

OpenStudy (jdoe0001):

well, let's leave the "12" on the side for now, is just a multiplier of the identity anyway

OpenStudy (jdoe0001):

1/6 + 1/6 = ?

OpenStudy (anonymous):

1/3

OpenStudy (jdoe0001):

thus \(\bf \left[sin \left(\cfrac{\pi}{6}\right)sin \left(\cfrac{\pi}{6}\right)\right] = \frac{1}{2}\left[ cos\left(0\right)-cos\left(\cfrac{2\pi}{6}\right)\right]\\ \implies \frac{1}{2}\left[ cos\left(0\right)-cos\left(\cfrac{\pi}{3}\right)\right]\\ \)

OpenStudy (jdoe0001):

so, what would be the cos(0)?

OpenStudy (anonymous):

pi/3 or 1/3?

OpenStudy (jdoe0001):

well, check your unit circle for the cos(0), or even your calculator :)

OpenStudy (anonymous):

oh its 1

OpenStudy (jdoe0001):

\(\bf \frac{1}{2}\left[ cos\left(0\right)-cos\left(\cfrac{\pi}{3}\right)\right] \implies \frac{1}{2}\left[ 1-cos\left(\cfrac{\pi}{3}\right)\right]\\ \cfrac{12\left[1-cos\left(\cfrac{\pi}{3}\right)\right]}{2}\)

OpenStudy (jdoe0001):

so now we bring in the multiplier, 12

OpenStudy (jdoe0001):

as you can pretty much see what the answer is :)

OpenStudy (jdoe0001):

\(\bf \cfrac{12\left[1-cos\left(\cfrac{\pi}{3}\right)\right]}{2}\\ \cfrac{\cancel{12}-\cancel{12}cos\left(\cfrac{\pi}{3}\right)}{\cancel{2}} \implies 6-6cos\left(\cfrac{\pi}{3}\right)\)

OpenStudy (anonymous):

ok thank you! i was working on it haha

OpenStudy (jdoe0001):

hehe

OpenStudy (jdoe0001):

yw